# KiDS-SBI: Simulation-Based Inference Analysis of KiDS-1000 Cosmic Shear

### v1

@@ 3.8 Variable depth
As alluded to in Sect. 3.6 and 3.7, when randomly sampling galaxies and their intrinsic ellipticities, the parameters which charaterise the probability distributions depend on the location on the sky, i.e. the pixel $m$ in question. Specifically, the number of galaxies which is sampled (i.e. ”observed”) within a given pixel depends on the average galaxy density, $n_{\mathrm{gal},m}^{(p)}$, measured within that pixel. The galaxy density varies from pixel to pixel as it depends on the observational depth of the survey in different parts of the sky, and on anisotropies and time variations in the atmospheric seeing (particularly, for surveys such as KiDS which only visit each pointing in the sky once at a given time). Since many parts of the sky are observed at different times, as is the case for ground-based telescopes such as the VLT Survey Telescope and the Visible and the Infrared Survey Telescope for Astronomy (VISTA) used for KiDS-1000, the atmospheric conditions and background light will be different for each pointing for a given exposure time. Thus, the depth of observations may change as seeing, the point spread functions and/or the background flux varies (for example, due to moonlight, zodiacal light or galactic absorption).
In addition, surveys such as KiDS often have overlapping pointings in order to ensure that the footprint is observed without gaps. However, this implies that galactic sources which happen to be located near the edge of a pointing will be observed more often in different overlapping fields than a source located in the centre of a pointing. Consequently, near the edge of a pointing, there will tend to be a higher signal-to-noise ratio which allows for deeper observations. Since these systematic effects in the observational depth can happen at fixed scales and/or have certain periodicities, they can induce significant systematic effects into galaxy clustering and weak gravitational lensing signals. For cosmic shear in KiDS-1000, the bias can be near $1\%$ in the signal and an average of ${\sim}10\%$ in the standard deviation of cosmic shear observables (Heydenreich et al., 2020; Joachimi et al., 2021).
@@
To take all these effects into consideration within the forward simulations, we calibrate the spatial variability of $n_{\mathrm{eff},m}^{(p)}$ and $\sigma_{\epsilon,m}^{(p)}$ from the KiDS-1000 measurements directly. Rather than defining maps over the entire survey footprint of both quantities for each tomographic bin, we use a map of a direct estimator of observational depth which also correlates with $n_{\mathrm{eff},m}^{(p)}$ and $\sigma_{\epsilon,m}^{(p)}$: the root-mean square of the background noise, $\sigma_{\mathrm{rms}}$. We found this quantity to be a good indicator of variable depth as it correlates well with the measured $n_{\mathrm{eff},m}^{(p)}$ and $\sigma_{\epsilon,m}^{(p)}$, while being uncorrelated with the measured photometric redshifts and magnitudes in KiDS-1000.
@@
As can be seen in Fig. 7 and  8, a map of $\sigma_{\mathrm{rms}}$ across the KiDS North and South fields clearly shows variations over different pointings as well as with areas of increased overlap. In addition, Fig. 9 shows that $n_{\mathrm{eff}}^{(p)}$ and $\sigma_{\epsilon}^{(p)}$ vary linearly with $\sigma_{\mathrm{rms}}$ for KiDS-1000 data. Thanks to this fact, we do not need to model the variation of $n_{\mathrm{eff},m}^{(p)}$ and $\sigma_{\epsilon,m}^{(p)}$ with the HEALPix pixel $m$ with individual maps for each tomographic bin $p$. Instead, using the linear relations shown in Fig. 9, we may assign a value for $n_{\mathrm{eff},m}^{(p)}$ and $\sigma_{\epsilon,m}^{(p)}$ for each $\sigma_{\mathrm{rms},m}$ (i.e. the mean value of level within each pixel $m$ contained within the footprint observed by KiDS-1000) given that
@@
where $a_{n_{\mathrm{eff}}}^{(p)}$ and $a_{\sigma_{\epsilon}}^{(p)}$ are the slopes of the linear fits shown in Fig. 9 of galaxy density and galaxy shape dispersion for a given tomographic bin, respectively, and $b_{n_{\mathrm{eff}}}^{(p)}$ and $b_{\sigma_{\epsilon}}^{(p)}$ are the associated y-intercepts. The values of these parameters are shown in Table 1. This drastically helps the performance of the forward simulations, as this allows to just use a single map of $\sigma_{\mathrm{rms}}$ to model the variable depth rather than requiring $2N_{\mathrm{tomo}}$ maps for it. Such a linear relation may not exist in other weak lensing surveys, but it should still be possible to numerically calibrate the relations dependence of $n_{\mathrm{eff}}^{(p)}$ and $\sigma_{\epsilon}^{(p)}$ on $\sigma_{\mathrm{rms}}$ to define an interpolation.
@@ 4 Measurements
Upon sampling galaxy shapes as well as shear, and applying all relevant instrumental systematics within KiDS-SBI, we measure the galaxy shapes, and the cosmic shear summary statistics thereof in the same manner as the measurements made from the KiDS-1000 gold sample’s weak lensing catalogue. The only difference between the measurements of the forward-modelled catalogues and the KiDS-1000 gold sample is the values of the galaxy weights, $w_{i}$. In the forward simulations, the galaxy weights are set to unity for all sampled galaxies, $w_{i}=1\forall i$, while the weights used in the measurements from the KiDS-1000 gold sample are given by the lensfit weights (see details in Sect. 2). To avoid biasing the statistics of the galaxy populations in the forward simulations with respect to the real data, we sample the galaxy positions and shapes from the galaxy densities, $n_{\mathrm{gal}}$, and intrinsic shape dispersions, $\sigma_{\epsilon}$, where the corrections for the lensfit weight and the multiplicative shear bias have been applied, as defined in Eqs. (18) and (19).
@@ 4.1 Shape measurements
where $w_{i}$ is the galaxy lensfit weight for galaxy $i$ in the KiDS-1000 gold sample (for the case of the simulations, $w_{i}=1\forall i$), $\langle\epsilon_{\mathrm{obs}}(\mathbf{\Theta})\rangle$ compensates for any additive biases (Asgari et al., 2019), i.e. the $c^{(p)}$ term and the mean of $\alpha^{(p)}$ in Eq. (27), and it is defined as the mean observed shape of all galaxies within tomographic bin $p$, and $\overline{M}^{(p)}$ is the mean multiplicative shear bias measured for each tomographic bin $p$ as shown in Fig. 5. Note that $\overline{M}^{(p)}$ is not necessarily the same value as the value of $M^{(p)}$ applied in Eq. (27). For each instance of the simulations, a different $M^{(p)}$ is drawn from a Gaussian probability distribution with mean $\overline{M}^{(p)}$ and the standard deviations shown in Fig. 5. Regardless of the value of $M^{(p)}$ drawn, the shear measurements of each simulation are corrected by the same mean $\overline{M}^{(p)}$. Any discrepancy between these values is going to introduce noise in the simulations which accounts for the uncertainty on the shear bias measurements.
@@ 4.2 Observed angular power spectra
We choose to compress the catalogues down to two-point statistics. The main reason for this is that the forward simulation pipeline is designed to be only percent-level accurate in two-point statistics. Higher-order correlations of the shear field are captured to a certain degree within the simulations, but their accuracy is limited (Hall & Taylor, 2022; Piras et al., 2023). In addition, the use of two-point statistics assures that the only distinguishing feature of our analysis when compared to the fiducial cosmic shear KiDS-1000 analyses (Loureiro et al., 2022; Asgari et al., 2021) is the dropping of the assumption of a Gaussian likelihood, so we can carry out a direct comparison between the analyses. This way we can ensure that any non-Gaussianities in our likelihood are either attributed to inherent non-Gaussianities in the likelihood of two-point statistics or non-Gaussianities induced by systematic effects.
Due to its computational efficiency when compared to spatial two-point correlation functions, we choose pseudo-Cls as our two-point statistic. We decompose the observed shear field, $\epsilon_{\mathrm{obs}}^{\mathrm{corr}\,(p)}(\theta_{m};\mathbf{\Theta})$, into its curl-free E-modes and its divergence-free B-modes as follows
@@
where $\mu\in\{1,2,3\}$ such that $1$ stands for the EE component, $2$ for the EB component and $3$ for the BB component. Note that $\tilde{C}_{\epsilon\epsilon}^{EB}(\ell;\mathbf{\Theta})=\tilde{C}_{\epsilon\epsilon}^{BE}(\ell;\mathbf{\Theta})$, since $\tilde{E}_{\ell m}\tilde{B}_{\ell m}^{*}(\mathbf{\Theta})=\tilde{B}_{\ell m}\tilde{E}_{\ell m}^{*}(\mathbf{\Theta})$.
@@
To account for the EE signal from the correlations between the intrinsic galaxy shapes, i.e. the shape noise bias, we subtract the mean shape noise power spectrum as follows
@@
Furthermore, we bin the noise-free observed angular power spectrum, $\tilde{C}^{(pq)}_{\epsilon\epsilon}(\ell;\mathbf{\Theta})$, into 8 log-spaced bins between $\ell=76$ and $\ell=1500$ in line with Loureiro et al. (2022). For this, we choose the pseudo-Cl binning scheme described in Brown et al. (2005) given by
@@
Note that we do not choose to deconvolve the pseudo-Cls to estimate the full-sky E-mode angular power spectra for cosmic shear. We find that with the sky-coverage of KiDS-1000 and the complexity of the geometry of the KiDS-1000 footprint, the mixing matrix (see Appendix D) is not necessarily invertible, and the deconvolution is not single-valued.
We therefore obtain for a given forward simulation a cosmic shear measurement like the one shown in Fig. 11 for a single run of the forward simulations. The calculation of the data vector calculated from the theory prediction shown in Fig. 11 is described in Appendix D. As can be seen from Figure 11, the measured cosmic shear signal is consistent with theory predictions which we find is the case at all cosmologies throughout the prior volume (see Appendix B). This shows that the log-normal random fields accurately recover the two-point statistics of the galaxy populations throughout parameter space, as expected.
As seen in Fig. 11, for the five tomographic bins in KiDS-1000, we obtain $28\times N_{\mathrm{tomo}}(N_{\mathrm{tomo}}+1)/$ 2 data points in $\tilde{C}^{(pq)}_{\epsilon\epsilon,L}(\mathbf{\Theta})$, i.e. a 120-dimensional data vector. The analysis choices described within Sect 3 and 4 in conjunction constitute the forward model assumed within this SBI analysis, and it is henceforth labelled as the anisotropic systematics model, where the plus sign is in reference to the fact that the model includes novel features such as variable depth, PSF variation, and dropping of the reduced shear approximation.
To summarise, each data vector depends on 12 parameters in $\mathbf{\Theta}$: 5 cosmological parameters ($\sigma_{8}$, the root-mean-square matter fluctuation over 8 Mpc$/h_{0}$; $\omega_{b}$, baryonic matter density; $\omega_{c}$, cold dark matter density; $n_{s}$, the scalar spectral index of the primordial density fluctuation power spectrum, $P_{\mathcal{R}}$; and $h_{0}$ [^14] Note that throughout we deviate from standard notation and denote the dimensionless Hubble constant as $h_{0}$., the normalised Hubble constant), and 7 astrophysical/nuisance parameters related to systematics: $A_{\mathrm{bary}}$, the baryonic feedback amplitude within the non-linear three-dimensional matter power spectrum; $A_{\mathrm{IA}}$, the galaxy intrinsic alignment amplitude of the NLA model; and five correlated $\delta_{z}$ parameters which define the shift in the mean of the source redshift distribution, $P(z|z_{\mathrm{ph}})$, of each tomographic bin. All other parameters on which the simulation depends are fixed from run to run (e.g. $\omega_{k}=0$, the equation of state of dark energy, $w=-1$, $\sum m_{\nu}=0.06$ eV, etc.). The exceptions to this are the seed used to sample the matter overdensities and the galaxies, and the amplitudes of the shear biases as described in Sect. 3.7. These values are varied from run to run in order to simulate cosmic variance, shape noise as well as the uncertainty in the shear bias, respectively.
@@ 5 Simulation-based inference (SBI)
To implement this, we create a bespoke interface that allows DELFI to be the sampler for CosmoSIS (Zuntz et al., 2015). Initially, the sampler chooses a fixed number of points within the considered parameter space based on the hypercube-generating algorithm from PyDOE [^16] https://github.com/tisimst/pyDOE with some minor modifications (see L23 for more details). Subsequently, as is outlined in Fig. 12, the measured pseudo-Cls for each evaluation of the parameters on the hypercube are compressed further using score compression (see Sect. 5.3). This reduces the dimensionality of the data vector to the size of the parameter vector (in this case, 7 dimensions). The compressed data is then used to train neural density estimators through DELFI (see Sect. 5.4).
As the ensemble of neural density estimators evaluates the sampling distribution throughout the hypercube, through the use of active learning, DELFI finds areas in parameter space where the likelihood is undersampled to select new parameter vectors (see Alsing et al. 2019 for details). As found in L23, this can halve the number of simulations needed to accurately train DELFI by drawing more simulations in the parameter region of interest when compared to a Latin hypercube alone. Once the ensemble of neural density estimators has reached the stopping criterion, controlled for over-fitting by comparison to a hold-out validation data set, we can calculate an effective likelihood from the learnt sampling distribution condition on the observed data vector (which can be a mock or measured from real data). We can then obtain a posterior distribution by sampling the product of the likelihood and the prior. This approach of learning the sampling distribution conditioned on the observed compressed data to yield an effective likelihood has the advantage that if the prior or data vector were to be changed, there is nothing new that must be re-trained within DELFI.

### v2

@@ 3.8 Variable depth
As alluded to in Sect. 3.6 and 3.7, when randomly sampling galaxies and their intrinsic ellipticities, the parameters which charaterise the probability distributions depend on the location on the sky, namely, the pixel $m$ in question. Specifically, the number of galaxies which is sampled (in other words, ’observed’) within a given pixel depends on the average galaxy density, $n_{\mathrm{gal},m}^{(p)}$, measured within that pixel. The galaxy density varies from pixel to pixel as it depends on the observational depth of the survey in different parts of the sky, and on anisotropies and time variations in the atmospheric seeing (particularly, for surveys such as KiDS which only visit each pointing in the sky once at a given time). Since many parts of the sky are observed at different times, as is the case for ground-based telescopes such as the VLT Survey Telescope and the Visible and the Infrared Survey Telescope for Astronomy (VISTA) used for KiDS-1000, the atmospheric conditions and background light will be different for each pointing for a given exposure time. Thus, the depth of observations may change as seeing, the point spread functions and/or the background flux varies (for example, due to moonlight, zodiacal light, or galactic absorption).
In addition, surveys such as KiDS often have overlapping pointings in order to ensure that the footprint is observed without gaps. However, this implies that galactic sources which happen to be located near the edge of a pointing will be observed more often in different overlapping fields than a source located in the centre of a pointing. Consequently, near the edge of a pointing, there will tend to be a higher S/N which allows for deeper observations. Since these systematic effects in the observational depth can happen at fixed scales and/or have certain periodicities, they can induce significant systematic effects into galaxy clustering and weak gravitational lensing signals. For cosmic shear in KiDS-1000, the bias can be near $1\%$ in the signal and an average of ${\sim}10\%$ in the standard deviation of cosmic shear observables (Heydenreich et al., 2020; Joachimi et al., 2021).
@@
To take all these effects into consideration within the forward simulations, we calibrate the spatial variability of $n_{\mathrm{eff},m}^{(p)}$ and $\sigma_{\epsilon,m}^{(p)}$ from the KiDS-1000 measurements directly. Rather than defining maps over the entire survey footprint of both quantities for each tomographic bin, we use a map of a direct estimator of observational depth which also correlates with $n_{\mathrm{eff},m}^{(p)}$ and $\sigma_{\epsilon,m}^{(p)}$: the root-mean square of the background noise, $\sigma_{\mathrm{rms}}$. This quantity gives the detection threshold of counts above the background, and is the dimensionless analogue to the detection threshold typically expressed in mag arcsec-2. Since it directly characterises the detection probability of an object as a function of its angular position, $\sigma_{\mathrm{rms}}$ incorporates variations in the magnitude limit or the seeing (Johnston et al., 2021). We found this quantity to be a good indicator of variable depth as it correlates well with the measured $n_{\mathrm{eff},m}^{(p)}$ and $\sigma_{\epsilon,m}^{(p)}$, while being uncorrelated with the measured photometric redshifts and magnitudes in KiDS-1000.
@@
As can be seen in Fig. 7 and  8, a map of $\sigma_{\mathrm{rms}}$ across the KiDS North and South fields clearly shows variations over different pointings as well as with areas of increased overlap. In addition, Fig. 9 shows that $n_{\mathrm{eff}}^{(p)}$ and $\sigma_{\epsilon}^{(p)}$ vary linearly with $\sigma_{\mathrm{rms}}$ for KiDS-1000 data. Thanks to this fact, we do not need to model the variation of $n_{\mathrm{eff},m}^{(p)}$ and $\sigma_{\epsilon,m}^{(p)}$ with the HEALPix pixel $m$ with individual maps for each tomographic bin $p$. Instead, using the linear relations shown in Fig. 9, we may assign a value for $n_{\mathrm{eff},m}^{(p)}$ and $\sigma_{\epsilon,m}^{(p)}$ for each $\sigma_{\mathrm{rms},m}$ (specifically the mean value of level within each pixel $m$ contained within the footprint observed by KiDS-1000) given that
@@
where $a_{n_{\mathrm{eff}}}^{(p)}$ and $a_{\sigma_{\epsilon}}^{(p)}$ are the slopes of the linear fits shown in Fig. 9 of galaxy density and galaxy shape dispersion for a given tomographic bin, respectively, and $b_{n_{\mathrm{eff}}}^{(p)}$ and $b_{\sigma_{\epsilon}}^{(p)}$ are the associated y-intercepts. The values of these parameters are shown in Table 13. This drastically helps the performance of the forward simulations, as this allows us to just use a single map of $\sigma_{\mathrm{rms}}$ to model the variable depth rather than requiring $2N_{\mathrm{tomo}}$ maps for it. Such a linear relation may not exist in other weak lensing surveys, but it should still be possible to numerically calibrate the relations dependence of $n_{\mathrm{eff}}^{(p)}$ and $\sigma_{\epsilon}^{(p)}$ on $\sigma_{\mathrm{rms}}$ to define an interpolation.
@@
We note that the galaxy selection function considered here does not account for systematic effects which have been previously corrected in KiDS data, such as source blending. Thanks to the selection and calibration of the data discussed in Sect. 2, the KiDS-1000 source catalogue as well as the simulations thereof can be treated as a sample of non-overlapping sources. Any residual effects from source blending which remain after the measurement and calibration stages of the data are included within the shear biases and weights. Consequently, when quantifying the variable depth of the survey, any residual effect on blending is compensated by the inclusion of the galaxy weights and the multiplicative shear bias in the definitions of the effective galaxy density and the intrinsic ellipticity dispersion of galaxies, in Eqs. (18) and (19).
@@ 4 Measurements
After sampling the galaxy shapes and shear, then applying all relevant instrumental systematics within KiDS-SBI, we can go on to measure the galaxy shapes, along with the cosmic shear summary statistics in the same manner as the measurements done for the KiDS-1000 gold sample’s weak lensing catalogue. The only difference between the measurements of the forward-modelled catalogues and the KiDS-1000 gold sample is the values of the galaxy weights, $w_{i}$. In the forward simulations, the galaxy weights are set to unity for all sampled galaxies, $w_{i}=1\forall i$, while the weights used in the measurements from the KiDS-1000 gold sample are given by the lensfit weights (see details in Sect. 2). To avoid biasing the statistics of the galaxy populations in the forward simulations with respect to the real data, we sample the galaxy positions and shapes from the galaxy densities, $n_{\mathrm{gal}}$, and intrinsic shape dispersions, $\sigma_{\epsilon}$, where the corrections for the lensfit weight and the multiplicative shear bias have been applied, as defined in Eqs. (18) and (19).
@@ 4.1 Shape measurements
where $w_{i}$ is the galaxy lensfit weight for galaxy $i$ in the KiDS-1000 gold sample (for the case of the simulations, $w_{i}=1\forall i$), $\langle\epsilon_{\mathrm{obs}}(\mathbf{\Theta})\rangle$ compensates for any additive biases (Asgari et al., 2019), that is the $c^{(p)}$ term and the mean of $\alpha^{(p)}$ in Eq. (27), and it is defined as the mean observed shape of all galaxies within tomographic bin $p$, and $\overline{M}^{(p)}$ is the mean multiplicative shear bias measured for each tomographic bin $p$ as shown in Fig. 5. We note that $\overline{M}^{(p)}$ is not necessarily the same value as the value of $M^{(p)}$ applied in Eq. (27). For each instance of the simulations, a different $M^{(p)}$ is drawn from a Gaussian probability distribution with mean $\overline{M}^{(p)}$ and the standard deviations shown in Fig. 5. Regardless of the value of $M^{(p)}$ drawn, the shear measurements of each simulation are corrected by the same mean $\overline{M}^{(p)}$. Any discrepancy between these values is going to introduce noise in the simulations which accounts for the uncertainty on the shear bias measurements.
@@ 4.2 Observed angular power spectra
We chose to compress the catalogues down to two-point statistics. The main reason for this is that the forward simulation pipeline is designed to be only percent-level accurate in two-point statistics. Higher-order correlations of the shear field are captured to a certain degree within the simulations, but their accuracy is limited (Hall & Taylor, 2022; Piras et al., 2023). In addition, the use of two-point statistics assures that the only distinguishing feature of our analysis when compared to the fiducial cosmic shear KiDS-1000 analyses (Loureiro et al., 2022; Asgari et al., 2021) is the dropping of the assumption of a Gaussian likelihood, so we can carry out a direct comparison between the analyses. This way we can ensure that any non-Gaussianities in our likelihood are either attributed to inherent non-Gaussianities in the likelihood of two-point statistics or non-Gaussianities induced by systematic effects.
Due to its computational efficiency when compared to spatial two-point correlation functions, we chose pseudo-Cls as our two-point statistic. We decompose the observed shear field, $\epsilon_{\mathrm{obs}}^{\mathrm{corr}\,(p)}(\theta_{m};\mathbf{\Theta})$, into its curl-free E-modes and its divergence-free B-modes as follows
@@
where $\mu\in\{1,2,3\}$ and $1$ stands for the EE component, $2$ for the EB component, and $3$ for the BB component. We note that $\tilde{C}_{\epsilon\epsilon}^{EB}(\ell;\mathbf{\Theta})=\tilde{C}_{\epsilon\epsilon}^{BE}(\ell;\mathbf{\Theta})$, since $\tilde{E}_{\ell m}\tilde{B}_{\ell m}^{*}(\mathbf{\Theta})=\tilde{B}_{\ell m}\tilde{E}_{\ell m}^{*}(\mathbf{\Theta})$.
@@
To account for the EE signal from the correlations between the intrinsic galaxy shapes, namely the shape noise bias, we subtract the mean shape noise power spectrum as follows
@@
Furthermore, we bin the noise-free observed angular power spectrum, $\tilde{C}^{(pq)}_{\epsilon\epsilon}(\ell;\mathbf{\Theta})$, into 8 log-spaced bins between $\ell=76$ and $\ell=1500$ in line with Loureiro et al. (2022). For this, we chose the pseudo-Cl binning scheme described in Brown et al. (2005) given by
@@
We note that we did not choose to deconvolve the pseudo-Cls to estimate the full-sky E-mode angular power spectra for cosmic shear. We find that with the sky-coverage of KiDS-1000 and the complexity of the geometry of the KiDS-1000 footprint, the mixing matrix (see Appendix D) is not necessarily invertible, and the deconvolution is not single-valued.
We therefore obtain, for a given forward simulation, a cosmic shear measurement like the one shown in Fig. 29 for a single run of the forward simulations. The calculation of the data vector calculated from the theory prediction shown in Fig. 29 is described in Appendix D. As can be seen from Fig. 29, the measured cosmic shear signal is consistent with theory predictions which we find is the case at all cosmologies throughout the prior volume (see Appendix B). This shows that the log-normal random fields accurately recover the two-point statistics of the galaxy populations throughout parameter space, as expected.
As seen in Fig. 29, for the five tomographic bins in KiDS-1000, we obtain $28\times N_{\mathrm{tomo}}(N_{\mathrm{tomo}}+1)/$ 2 data points in $\tilde{C}^{(pq)}_{\epsilon\epsilon,L}(\mathbf{\Theta})$, namely, a 120D data vector. The analysis choices described within Sect 3 and 4 in conjunction constitute the forward model assumed within this SBI analysis, and it is henceforth labelled as the anisotropic systematics model, in reference to the fact that the model includes novel features such as variable depth, PSF variation, and dropping of the reduced shear approximation.
To summarise, each data vector depends on 12 parameters in $\mathbf{\Theta}$: 5 cosmological parameters ($\sigma_{8}$, the root-mean-square matter fluctuation over 8 Mpc$/h_{0}$; $\omega_{b}$, baryonic matter density; $\omega_{c}$, cold dark matter density; $n_{s}$, the scalar spectral index of the primordial density fluctuation power spectrum, $P_{\mathcal{R}}$; and $h_{0}$ [^15] Note: throughout we deviate from standard notation and denote the dimensionless Hubble constant as $h_{0}$., the normalised Hubble constant), and 7 astrophysical and nuisance parameters related to systematics: $A_{\mathrm{bary}}$, the baryonic feedback amplitude within the non-linear 3D matter power spectrum; $A_{\mathrm{IA}}$, the galaxy intrinsic alignment amplitude of the NLA model; and five correlated $\delta_{z}$ parameters which define the shift in the mean of the source redshift distribution, $P(z|z_{\mathrm{ph}})$, of each tomographic bin. All other parameters on which the simulation depends are fixed from run to run (such as $\omega_{k}=0$, the equation of state of dark energy, $w=-1$, $\sum m_{\nu}=0.06$ eV, etc.). The exceptions to this are the seed used to sample the matter overdensities and the galaxies, and the amplitudes of the shear biases as described in Sect. 3.7. These values are varied from run to run in order to simulate cosmic variance, shape noise as well as the uncertainty in the shear bias, respectively.
@@ 5 Simulation-based inference (SBI)
To implement this, we create a bespoke interface that allows DELFI to be the sampler for CosmoSIS (Zuntz et al., 2015). Initially, the sampler chooses a fixed number of points within the considered parameter space based on the hypercube-generating algorithm from PyDOE [^17] https://github.com/tisimst/pyDOE with some minor modifications (see L23 for more details). Subsequently, as is outlined in Fig. 11, the measured pseudo-Cls for each evaluation of the parameters on the hypercube are compressed further using score compression (see Sect. 5.3). This reduces the dimensionality of the data vector to the size of the parameter vector (in this case, seven dimensions). The compressed data is then used to train neural density estimators through DELFI (see Sect. 5.4).
As the ensemble of neural density estimators evaluates the sampling distribution throughout the hypercube, through the use of active learning, DELFI finds areas in parameter space where the likelihood is undersampled to select new parameter vectors (see Alsing et al. 2019 for details). As found in L23, this can halve the number of simulations needed to accurately train DELFI by drawing more simulations in the parameter region of interest when compared to a Latin hypercube alone. Once the ensemble of neural density estimators has reached the stopping criterion, controlled for over-fitting by comparison to a hold-out validation data set, we can calculate an effective likelihood from the learnt sampling distribution conditioned on the observed data vector (which can be a mock or measured from real data). We can then obtain a posterior distribution by sampling the product of the likelihood and the prior. This approach of learning the sampling distribution conditioned on the observed compressed data to yield an effective likelihood has the advantage that if the prior or data vector were to be changed, there is nothing new that must be re-trained within DELFI.